Suppose that when converting to the same currency values, the nominal GDP per capita in the fictional country of Islandia is 25 percent higher than the nominal GDP per capita in the fictional country of Mountainia. However, the purchasing power for the same amount of Islandia currency is about 40 percent lower in Islandia than in Mountainia. If we use Islandia as the base country for comparison, the PPP-adjusted GDP per capita in Mountainia i(Click to select) ts nominal GDP. less than Which country has the higher average standard-of-living?greater thar equal to Mountainia Both are equal OIt is not possible to determine which is higher O IslandiaPurchasing power parity (PPP): is a reason why all economies have left the gold standard. is as commonly accepted as the law of demand. O almost never holds completely. represents the universality of exchange rate systems.
Answer #1

Assume

NGDP per capita of Mountainia = X

NGDP percapita of Islandia= X (1+25%) = 1.25X

Purchasing Power of Mountainia = Y

Purchasing power of Islandia = Y(1-40%) = 0.6Y

Which says that currency value of Islandia is less than currency value of Mountainia. So when Islandia is considered as base country than adjusted GDP of Mountainia is less than its Nominal GDP as the currency value of Islandia is low than Mountainia.

Answer: Is less than

2) Answer: Islandia

As the GDP per capita of Islandia is more than Mountainia it shows that Islandia is having a higher growth rate when compared to Mountainia which means that people of Islandia have higher average standard of living.

3) Answer: represents the universality of exchange rate system

PPP is the alternative to exchange rate system which is used to compare different countries purchasing power. It follows the rule of law of one price.